OpenStudy (anonymous):

A quadratic equation with real coefficients and leading coefficient 1, has x = -bi as a root. Write the equation in general form.

OpenStudy (mathmate):

A polynomial with real coefficients has complex roots in conjugate pairs. Hence if one root is -bi, the conjugate root is +bi. The quadratic is then f(x)=a(x+bi)(x-bi). If the leading coefficient (coefficient of the term of highest degree, namely x^2) is one, then a=1, giving f(x)=(x+bi)(x-bi), Expand the expression and write in general form to complete the solution.

OpenStudy (anonymous):

how will i write in in general form? :)

OpenStudy (anonymous):

\[x ^{2}-bi^2 =0 ???

OpenStudy (mathmate):

Almost! But do not forget \(ab^2\) does not equal \((ab)^2=a^2b^2\), and that \(i^2=-1\) \(x ^{2}-(bi)^2 =0\) Expand the second term and simplify to get the general form.